# Communication-Efficient and Tensorized Federated Fine-Tuning of Large Language Models

### v1

@@ 6 Conclusion
This work presents a novel federated and tensorized fine-tuning framework, addressing the key challenges of communication overhead and data heterogeneity in FL. FedTT leverages tensorized adapters to significantly reduce communication costs while maintaining high performance, achieving up to a 10×10$\times$10 × reduction in trainable parameters compared to existing methods. Additionally, FedTT+ enhances robustness in cross-silo FL by adaptively freezing portions of tensor factors, further optimizing parameter efficiency. Extensive experiments with models like BERT and LLaMA-2 demonstrate the effectiveness of FedTT and FedTT+ in both cross-silo and large-scale cross-device FL settings, offering a scalable and communication-efficient solution for fine-tuning large language models in distributed environments.
@@
Although the primary motivation for using FL is to enhance data privacy, different algorithms provide varying levels of privacy protection, as privacy leakage is inevitable in most methods. A key experiment missing from our work is a comparison of our algorithm with others in terms of privacy preservation. Given that a lower number of trainable parameters could intuitively lead to better privacy in FL settings, our framework is likely to offer strong privacy-preserving capabilities, which presents an interesting area for future investigation.
@@ A.1 Dataset Setup
We began our experiments using the Generalized Language Understanding Evaluation (GLUE) benchmark [48], which includes a range of natural language understanding tasks. These tasks consist of perceptual analysis (SST2 [50]), similarity and paraphrase tasks (MRPC, QQP [51]), and natural language reasoning (MNLI, QNLI [52, 53]). The utilized metrics for the GLUE benchmark are summarized in Table 4. We recorded the best validation results after $100$ communication rounds for cross-silo FL and after $1000$ communication rounds for large-scale cross-device FL. For the DeBERTa-Base models on the QQP task, we randomly selected $1000$ samples from the validation set and reported the highest accuracy achieved.
@@
We then selected two multiple-choice tasks (COPA and ReCoRD) from the SuperGLUE benchmark [54] and a question-answering generation task (SQuAD [55]). The metrics used for evaluation are summarized in Table 5.
@@ A.3 Additional Details for Table 1
RoBERTa-Large (FedTT+) $27K$ $\bm{27K}$ $\bm{27K}$ $\bm{27K}$ $\bm{27K}$
Fed. Set. Model & Method # Param. COPA ReCoRD SQuAD Avg.
large-scale FL LLaMA2-7B (LoRAr=8) $4.19M$ $87$ $81.0$ $90.49$ $86.16$
LLaMA2-7B (FedTT) $0.52M$ $90$ $80.1$ $89.32$ $86.47$
cross-silo FL LLaMA2-13B (LoRAr=8) $6.55M$ $89$ $83.6$ $91.07$ $87.89$
LLaMA2-13B (FedTT) $0.64M$ $89$ $83.5$ $90.06$ $87.52$
LLaMA2-13B (FedTT+) $0.18M$ $88$ $83.1$ $89.12$ $86.74$
@@ A.4 Additional Details for Table 2
We follow the experimental setup of [14], using RoBERTa-Base models with the GLUE dataset in both cross-silo and large-scale cross-device FL configurations. The learning rate was selected from $[1e-2,5e-3,1e-3,5e-4,1e-4,5e-5]$, and batch size from $[16,32]$ across all tasks and methods. For the MRPC and SST-2 tasks, we used a cross-silo FL setup with $10$ clients, while for the other tasks, we employed a large-scale cross-device FL setup with $1000$ clients, selecting $10$ clients per round. Accuracy results for Prefix, Adapter, LoRA, and BitFit methods are sourced from [14]. In addition to reporting the accuracy of the FedTT method, we also included the P-Tuning method with prompt lengths $[128,768]$, and IA3 [56] method, to further enrich the experiments.
@@ A.5 Additional Details for Table 3
We adopt the same settings as [37], using RoBERTa-Large models with the GLUE dataset in cross-silo FL scenarios. To simulate varying levels of data heterogeneity, we report the accuracy of our method for $3$ clients (i.i.d. data distribution), $20$ clients (moderate heterogeneity), and $50$ clients (high heterogeneity), with $20$ local updates. The learning rate is set to $31e-$ 3 for all tasks, with a batch size of $64$ for SST-2 and $32$ for other tasks. In [37], to reduce the number of trainable parameters, LoRA is applied selectively to certain layers: the last $6$ layers for SST-2, and the last $9$ layers for QNLI, QQP, and MNLI. Consequently, the number of trainable parameters varies depending on the task. We provide a comparison of the number of trainable parameters for LoRA, FFA-LoRA, RoLoRA, and our proposed FedTT+ method in Table 8. Accuracy results for LoRA, FFA-LoRA, and RoLoRA are sourced from [37].
Appendix B Performance Comparison on the Large-Scale Language Models
Due to the large number of parameters in LLaMA-2, it is rarely used in FL scenarios. None of the existing works on LLMs provide results for such models. However, by employing FedTT, the significantly reduced number of communication overhead allows us to utilize these models. We use SuperGLUE tasks [54] and a generation task (SQuAD [55]) to compare FedTT and FedTT+ with LoRA. The results are shown in Table 9. We used the LLaMA2-7B model to simulate large-scale cross-silo FL with $1000$ clients, where $10$ of them are chosen randomly in each round for training. FedTT has performance nearly identical to LoRA with a rank of $8$, while achieving about 10×10$\times$10 × lower communication overhead. For cross-silo FL, we utilized the LLaMA2-13B model with $10$ clients. As seen in Table 9, both FedTT and FedTT+ demonstrate performance close to LoRA, with 10×10$\times$10 × and 30×30$\times$30 × lower communication overhead, respectively.

### v2

@@ 6 Conclusion
This work presents a novel federated and tensorized fine-tuning framework, addressing the key challenges of communication overhead and data heterogeneity in FL. FedTT leverages tensorized adapters to significantly reduce communication costs while maintaining high performance, achieving up to a 10×10$\times$10 × reduction in trainable parameters compared to existing methods. Additionally, FedTT+ enhances robustness in cross-silo FL by adaptively freezing portions of tensor factors, further optimizing parameter efficiency. Extensive experiments with models like BERT and LLaMA-2 demonstrate the effectiveness of FedTT and FedTT+ in both cross-silo and LSCD FL settings, offering a scalable and communication-efficient solution for fine-tuning LLMs in distributed environments.
@@
Acknowledgments
This work was supported by the National Science Foundation under Grant 2419982 and Grant 2342253. Z. Zhang and Y. Yang were supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Artificial Intelligence for Science program, under contract DE-SC0025390. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0030039.
@@ A.1 Dataset Setup
We begin our experiments using the Generalized Language Understanding Evaluation (GLUE) benchmark Wang et al. (2018), which includes a range of natural language understanding tasks. These tasks consist of perceptual analysis (SST2 Socher et al. (2013)), similarity and paraphrase tasks (MRPC, QQP Dagan et al. (2005)), and natural language reasoning (MNLI, QNLI Williams et al. (2017); Rajpurkar et al. (2018)). The utilized metrics for the GLUE benchmark are summarized in Table 8. We record the best validation results after $100$ communication rounds for cross-silo FL and after $1000$ communication rounds for large-scale cross-device FL. For the DeBERTa-Base models on the QQP task, we randomly select $1000$ samples from the validation set and report the highest accuracy achieved.
@@
We then select two multiple-choice tasks (COPA and ReCoRD) from the SuperGLUE benchmark Wang et al. (2019) and a question-answering generation task (SQuAD Rajpurkar et al. (2016)). The metrics used for evaluation are summarized in Table 9.
@@ A.3 Additional Details for Table 1
RoBERTa-Large (FedTT) $39K$ $55K$ $55K$ $55K$ $51K$
RoBERTa-Large (FedTT+) $28K$ $\bm{28K}$ $\bm{28K}$ $\bm{28K}$ $\bm{28K}$
@@ A.4 Additional Details for Table 2
We follow the experimental setup of Zhang et al. (2023), using RoBERTa-Base models with the GLUE dataset in both cross-silo and large-scale cross-device FL configurations. The learning rate was selected from $[1e-2,5e-3,1e-3,5e-4,1e-4,5e-5]$, and batch size from $[16,32]$ across all tasks and methods. For the MRPC and SST-2 tasks, we use a cross-silo FL setup with $10$ clients, while for the other tasks, we employ a large-scale cross-device FL setup with $1000$ clients, selecting $10$ clients per round. Accuracy results for Prefix, Adapter, LoRA, and BitFit methods are sourced from Zhang et al. (2023). In addition to reporting the accuracy of the FedTT method, we also includ the P-Tuning method with prompt lengths $[128,768]$, and IA3 Liu et al. (2022) method, to further enrich the experiments.
@@ A.5 Additional Details for Table 3
We adopt the same settings as Chen et al. , using RoBERTa-Large models with the GLUE dataset in cross-silo FL scenarios. To simulate varying levels of data heterogeneity, we report the accuracy of our method for $3$ clients (i.i.d. data distribution), $20$ clients (moderate heterogeneity), and $50$ clients (high heterogeneity), with $20$ local updates. The learning rate is set to $31e-$ 3 for all tasks, with a batch size of $64$ for SST-2 and $32$ for other tasks. To ensure a comparable number of trainable parameters, we make the last $6$ layers trainable for the SST-2 dataset and the last $9$ layers for the other datasets in LoRA, FFA-LoRA, RoLoRA, and FedTT. In contrast, for FedTT+, all layers are made trainable. We provide a comparison of the number of trainable parameters for LoRA, FFA-LoRA, RoLoRA, and our proposed FedTT, and FedTT+ methods in Table 12. Accuracy results for LoRA, FFA-LoRA, and RoLoRA are sourced from Chen et al. .
A.6 Additional Details for Table 5
Due to the large number of parameters in LLaMA-2, it is rarely used in FL scenarios, and few existing works provide results for such models. However, FedTT significantly reduces communication overhead, making it feasible to utilize these models in FL settings. We evaluate FedTT and FedTT+ against LoRA using SuperGLUE tasks Wang et al. (2019) and a generation task (SQuAD Rajpurkar et al. (2016)). The results are presented in Table 5.
For each task, we randomly select $1000$ samples for training and $1000$ for validation, reporting the best validation accuracy. We use a learning rate of $41e-$ 4, batch size of $2$, and $3$ local updates across all tasks. In large-scale cross-silo FL, we simulate training with the LLaMA2-7B model across $1000$ clients, selecting $10$ randomly per round. FedTT achieves performance nearly identical to LoRA with a rank of $8$, while reducing communication overhead by approximately 10×10$\times$10 ×.
For cross-silo FL, we utilize the LLaMA2-13B model with $10$ clients. As shown in Table 5, both FedTT and FedTT+ achieve performance close to LoRA while reducing communication overhead by approximately 10×10$\times$10 × and 30×30$\times$30 ×, respectively.
Appendix B Additional Experiment on Data Heterogeneity
Table 13: Comparison of cross-silo FL methods using RoBERTa-Large models under varying degrees of data heterogeneity. The accuracies for LoRA and FFA-LoRA methods are sourced from Sun et al. (2024).
Data Dist. Model & Method # Param. SST-2 QNLI QQP MNLI Avg.
i.i.d. RoBERTa-Large (LoRAr=8) $1.57M$ $94.42$ $91.38$ $84.47$ $86.90$ $89.29$
RoBERTa-Large (FFA-LoRAr=8) $0.79M$ $95.14$ $92.64$ $86.31$ $87.13$ $90.30$
RoBERTa-Large (FedTT+) $\bm{0.03M}$ $\bm{95.41}$ $\bm{94.05}$ $\bm{88.15}$ $\bm{88.65}$ $\bm{91.56}$
mild het. RoBERTa-Large (LoRAr=8) $1.57M$ $93.55$ $91.36$ $84.41$ $87.01$ $89.08$
RoBERTa-Large (FFA-LoRAr=8) $0.79M$ $94.10$ $91.62$ $85.33$ $87.04$ $89.52$
RoBERTa-Large (FedTT+) $\bm{0.03M}$ $\bm{95.53}$ $\bm{93.54}$ $\bm{87.91}$ $\bm{88.45}$ $\bm{91.36}$
sever het. RoBERTa-Large (LoRAr=8) $1.57M$ $94.32$ $88.95$ $83.51$ $82.03$ $87.20$
RoBERTa-Large (FFA-LoRAr=8) $0.79M$ $94.32$ $90.35$ $84.35$ $85.05$ $88.52$
RoBERTa-Large (FedTT+) $\bm{0.03M}$ $\bm{95.64}$ $\bm{91.67}$ $\bm{86.66}$ $\bm{87.73}$ $\bm{90.42}$
Data heterogeneity is a common challenge in most practical scenarios Hajihashemi and Shen . In FL, local models on individual clients can diverge from the global model’s optimal state, resulting in slower convergence Hsieh et al. (2020); Li et al. (2020). This issue is particularly pronounced when training LLMs in federated environments, as data heterogeneity can severely impact model performance Zhang et al. (2023).
As mentioned before, three primary factors contribute to data heterogeneity: (1) the number of clients, (2) the number of local updates, and (3) non-i.i.d. data distribution. In Section 5.3 of the paper, which evaluates the heterogeneity robustness of FedTT+, we considered the effects of the number of clients and local updates. Specifically, we increased the number of clients to intensify heterogeneity. In this section, we conduct an additional experiment to evaluate the performance of FedTT+ under varying levels of data heterogeneity. Similar to Sun et al. (2024), we train the RoBERTa-Large model and consider three levels of data distribution for three clients:
  • i.i.d. Data Distribution: Data is evenly distributed across all clients.
  • Mild Heterogeneity: For binary classification tasks, data is split as $[0.15,0.85]$, $[0.85,0.15]$, and $[0.5,0.5]$ among three clients. For three-class classification tasks, the splits are $[0.6,0.2,0.2]$, $[0.2,0.6,0.2]$, and $[0.2,0.2,0.6]$.
  • Severe Heterogeneity: For binary classification tasks, data is split as $[0.05,0.95]$, $[0.95,0.05]$, and $[0.5,0.5]$. For three-class classification tasks, the splits are $[0.9,0.05,0.05]$, $[0.05,0.9,0.05]$, and $[0.05,0.05,0.9]$.
To further amplify heterogeneity, we used $10$ local updates. We followed the exact experimental setup described in Sun et al. (2024), allowing us to directly compare our results with those reported for LoRA and FFA-LoRA. The results are presented in Table 13. As shown in the Table, FedTT+ consistently outperforms LoRA and FFA-LoRA across different heterogeneity settings while utilizing fewer parameters.
Table 14: Communication cost analysis for Table 1.
Method Up-link Message Size (KB) # Communication Round Total transmitted messages (KB)
SST-2 QNLI QQP MNLI SST-2 QNLI QQP MNLI
LoRAr=4 $586$ $2$ $2$ $2$ $2$ $1172$ $1172$ $1172$ $1172$
RoLoRAr=4 $312$ $2$ $2$ $1$ $2$ $624$ $624$ $312$ $624$
FedTT $234$ $2$ $2$ $2$ $2$ $468$ $\bm{468}$ $468$ $468$
FedTT+ $\bm{78}$ $5$ $6$ $3$ $3$ $\bm{390}$ $\bm{468}$ $\bm{234}$ $\bm{234}$
Table 15: Communication cost analysis for sever heterogeneity in Table 3.
Method Up-link Message Size (KB) # Communication Round Total transmitted messages (KB)
SST-2 QNLI QQP MNLI SST-2 QNLI QQP MNLI
RoLoRAr=2 $133$ $3$ $10$ $11$ $11$ $399$ $1330$ $1463$ $1463$
FedTT $199$ $3$ $3$ $3$ $3$ $597$ $597$ $597$ $597$
FedTT+ $\bm{109}$ $3$ $3$ $3$ $3$ $\bm{327}$ $\bm{327}$ $\bm{327}$ $\bm{327}$
Appendix C Communication Cost Analysis
In this section, we provide communication cost analysis for Table 1 and Table 3. Specifically, we compare the proposed method with baselines in terms of communication cost, following Guo et al. (2024). The analysis includes:
  • Up-link message size (in KB) for each communication round.
  • Number of communication rounds needed to reach the predefined target performance on the SST-2, QNLI, QQP, and MNLI datasets.
  • Total transmitted messages (in KB).
The target performance is defined as $95$% of the prediction accuracy reported in Table 1 and Table 3.
In a federated learning system, two key parameters influence communication efficiency: the up-link message size and the total size of transmitted messages, computed as the product of the number of communication rounds and the up-link message size. We provide detailed comparisons in Table 14 and Table 15. As demonstrated, our proposed algorithms effectively minimize both the up-link message size and the total number of transmitted messages.
Appendix D Privacy
While LLMs excel in performance due to their transformer-based architecture and vast parameter count, their ability to memorize and inadvertently reveal sensitive information from training data is a concern Carlini et al. (2021); Rathod et al. (2025). A widely adopted framework for mitigating such privacy risks is DP Dwork et al. (2006), which provides formal guarantees against data leakage.
In this section, we first outline key definitions of DP and introduce the DP-SGD algorithm. We then establish a theoretical privacy guarantee within the DP framework.
Definition 1. ($(\epsilon,\delta)$-Differential Privacy Dwork et al. (2006)]) A randomized mechanism $\mathcal{M}:\mathcal{D}\to\mathcal{R}$ satisfies $(\epsilon,\delta)$-differential privacy if for any two adjacent datasets $D,D^{\prime}\in\mathcal{D}$ that differ in at most one data point, and for any subset of possible outputs $S\subseteq\mathcal{R}$, the following holds:
$\Pr[\mathcal{M}(D)\in S]\leq e^{\epsilon}\Pr[\mathcal{M}(D^{\prime})\in S]+\delta.$
DP ensures that the mechanism $\mathcal{M}$ provides privacy guarantees by limiting the impact of any single data point on the output, with $\epsilon$ controlling the privacy loss and $\delta$ allowing for a small probability of failure in the guarantee.
Differentially Private Stochastic Gradient Descent (DP-SGD) Algorithm Abadi et al. (2016): DP-SGD is a modification of the SGD algorithm designed to provide differential privacy guarantees. It achieves this by introducing two key modifications:
  • Gradient Clipping: To ensure that individual data points do not have a disproportionate influence on the model update, each per-sample gradient is clipped to a fixed norm $C$.
  • Noise Addition: After aggregating the clipped gradients, Gaussian noise $z\sim\mathcal{N}\left(0,C^{2}\sigma^{2}I\right)$ is added to the sum of clipped gradients in a batch $\mathcal{B}$ from the dataset $\mathcal{D}$ before updating the model parameters. This helps obscure the contribution of any single data point.
The noisy sum of clipped gradients is computed as:
$\bar{g}=\frac{\sum_{i\in\mathcal{B}}\operatorname{Clip}(\nabla f_{i},C)+z}{|\mathcal{B}|},$
which is then used to update the model. Here, the noise scale $\sigma$ is determined based on sequential composition rules, given the privacy parameters $\epsilon$ and $\delta$, the number of iterations $T$, and the sampling probability $q=|\mathcal{B}|/|\mathcal{D}|$.
Although other techniques like secure aggregation have been explored in privacy-preserving FL Behnia et al. (2024), our focus here is on DP. In federated learning with DP, the level of privacy protection depends on whether the aggregation server is trusted by the clients. In global DP, the server is trusted, allowing clients to send raw model updates, while the server applies differential privacy by adding noise to the aggregated updates before releasing them McMahan et al. (2017b). In contrast, local DP assumes an untrusted server, requiring each client to add noise to their updates before transmission, ensuring privacy even if the server is compromised Wu et al. (2020). In Section 5.6, we adopt the stronger local DP approach to guarantee robust privacy protection without relying on a trusted server.
Proposition 1. The mechanism for updating FedTT using locally run DP-SGD and FedAvg satisfies $(\epsilon,\delta)$-DP, given that $\forall i$, the sampling probability is defined as $q=\left|\mathcal{B}_{i}\right|/\left|\mathcal{D}_{i}\right|$. Furthermore, the total number of local updates per client is denoted as $T$, and $\sigma$ is chosen as
$\sigma=O\left(\frac{q\sqrt{T\log(1/\delta)}}{\epsilon}\right).$
Proof. Similar to Sun et al. (2024).